On heating 0.220g of a metallic oxide in hydrgen,0.045g of water is formed.If the equivalent weight of the metal is E, then what is the value of E/9?
 

Grade:11

1 Answers

Bhavya
7 years ago
Dear student
Let the metallic oxide be MO
MO + H2 -----> H2O + M
Now, by law of equivalence i know that,
equivalence of MO = equivalence of H2O
weight of MO/Equivalent weight of MO = weight of H2O/equivalence of H2O
0.22/ E + 8 = 0.045 / 9
On solving you will get,
E/9 = 4